OpenStudy (anonymous):

Find the total price using each unit price: 5ft at $3 per foot?

terenzreignz (terenzreignz):

3 dollars per foot and you have 5 feet... how much for the whole thing, then?

OpenStudy (anonymous):

5*3=15 so... 5ft*3$=15$

OpenStudy (anonymous):

Here @BoraBora6636, I think this might be a little easier to digest: \[5 ft \times \frac{ $3 }{ ft }= $15\]